# Vulnerability Webs:

### v2

@@ Appendix B Exploiting sparse matrix operations for updating variational parameters
$\Lambda_{ik}$ $\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)=$ $\left(\sum_{\bm{x}\in\mathcal{X}}\sum_{d\in\{0,1\}}\bm{\Gamma}(d)\circ\bm{\Delta}(\bm{x})\,\bm{\Xi}^{(s)}\,\bm{\Pi}^{(s)}\left(d,\bm{x},\bm{\Xi}^{(s)}\right)+\left(\bm{\Gamma}(d)\circ\bm{\Delta}(\bm{x})\right)^{\top}\,\bm{\Xi}^{(s)}\,\bm{\Pi}^{(s)}\left(d,\bm{x},\bm{\Xi}^{(s)}\right)^{\top}\right)_{ik}.$
@@
$\bm{\pi}^{(s+1)}(1,\bm{x})=$ $\big{(}\left(\bm{\Xi}^{(s+1)}\right)^{\top}\bm{g}\circ\bm{X_{1}}^{x_{1}}\circ(\bm{J}-\bm{X_{1}})^{1-x_{1}}\circ\cdots\circ\bm{X_{p}}^{x_{p}}\circ(\bm{J}-\bm{X_{p}})^{1-x_{p}}\left(\bm{\Xi}^{(s+1)}\right)\big{)}\oslash$
$\left(\left(\bm{\Xi}^{(s+1)}\right)^{\top}\bm{X_{1}}^{x_{1}}\circ(\bm{J}-\bm{X_{1}})^{1-x_{1}}\circ\cdots\circ\bm{X_{p}}^{x_{p}}\circ(\bm{J}-\bm{X_{p}})^{1-x_{p}}\left(\bm{\Xi}^{(s+1)}\right)\right),$
@@
enabling the evaluation of $\bm{\Pi}\left(d,\bm{x},\bm{\Xi}^{(s)}\right)$ for $\bm{x}\in\mathcal{X}$. Following Dahbura et al. (2021), we can still evaluate this matrix by sparse matrix operations.
@@
In sum, we have shown how to compute the terms $\Omega_{ik}\left(\bm{0},\bm{0},\bm{\Xi}^{(s)}\right)$ and $\Lambda_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(s)}\right)$ through matrix multiplications. Plugging in these terms into (B) enables a fast computation of $\Omega_{ik}(\bm{g},\bm{x},\bm{\Xi}^{(s)})$, which is the computational bottleneck for evaluating the quadratic term needed to update $\bm{\Xi}^{(s)}$ to $\bm{\Xi}^{(s+1)}$.
B.1 STEP 2: Estimation of structural utility parameters
In the second step, we condition on the approximate block structure estimated in the first step $\widehat{\bm{z}}$ and estimate the structural payoff parameters. We compute the conditional probability of a link within types and between types
$p_{ij}(\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\gamma;\widehat{\bm{z}})$ $=$ $\begin{cases}\Lambda\left(u_{ij}(\alpha_{w},\bm{\beta}_{w})+u_{ji}(\alpha_{w},\bm{\beta}_{w})+4\gamma\sum_{r\neq i,j}I_{ijr}g_{jr}g_{ir}\right)&\text{ if }\widehat{\bm{z}}_{i}=\widehat{\bm{z}}_{j}\\ \Lambda\left(u_{ij}(\bm{\alpha},\bm{\beta})+u_{ji}(\bm{\alpha},\bm{\beta})\right)&\text{ otherwise }\end{cases}$
where $I_{ijr}=1$ if $\widehat{\bm{z}}_{i}=\widehat{\bm{z}}_{j}=\widehat{\bm{z}}_{r}$ and $I_{ijr}=0$ otherwise; and $\Lambda(u)=e^{u}/(1+e^{u})$ is the logistic function.
The maximum pseudolikelihood estimator (MPLE) solves the following maximization problem
$(\widehat{\bm{\alpha}},\widehat{\bm{\beta}},\widehat{\gamma})$ $=$ $\arg\max_{\bm{\alpha},\bm{\beta},\gamma}\ell_{PL}(\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\gamma;\widehat{\bm{z}})$
$=$ $\arg\max_{\bm{\alpha},\bm{\beta},\gamma}\;\sum_{i\neq j}^{n}\,\left(g_{ij}\log p_{ij}(\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\gamma;\widehat{\bm{z}})+(1-g_{ij})\log(1-p_{ij}(\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\gamma;\widehat{\bm{z}}))\right)$
In practice the estimator maximizes the log of the product of conditional probabilities of linking. The asymptotic theory for this estimator is in Boucher and Mourifie (2017) and Stewart and Schweinberger (2023). It can be shown that the estimator is consistent and asymptotically normal. As long as the estimator for $\bm{z}$ provides consistent estimates, the estimator for the structural parameters is well-behaved.
@@ B.1 STEP 2: Estimation of structural utility parameters
Appendix C Additional figures and estimates
C.1 Comparison with other models
@@
In Figure A1 we show the simulated statistics generated by a model with $K=1$. In aggregate, this model performs very poorly. This suggests that the unobserved heterogeneity can help fit the model better.
Figure A1: Model fit, ERGM ($K=1$)
@@
We compare our model with $K=10$ unobserved types to a stochastic blockmodel with $K=10$ unobserved types. In essence this is the same as our model, but we exclude the externality. As expected the stochastic blockmodel does a worse job predicting the externality within blocks, as shown in Figure A2.
Figure A2: Model fit, stochastic blockmodel with $K=10$
@@
We also did many robustness checks to choose the number of types $K$. In Figures A3 and A4 we report the boxplot for the fit of models with $K=5$ and $K=15$, respectively.
Figure A3: Model fit, model with $K=5$
Figure A4: Model fit, model with $K=15$
@@
From the graph we can conclude that increasing or decreasing the number of types does not seem to help in estimating the externality.
@@
Within Between
Estimate Estimate
(Std. Error) (Std. Error)
Edges ($\alpha_{1}$) $-5.703$ $-7.979$
(0.274) (0.006)
@@
Maturity ($\beta_{1}$) 1.767 $-0.222$
(0.303) (0.011)
$\log(n)\times$ Maturity ($\beta_{1}$) $-0.12$
Popularity ($\beta_{1}$) 1.071 $-0.07$
(0.308) (0.011)
$\log(n)\times$Popularity ($\beta_{1}$) $-0.117$
Size ($\beta_{1}$) 1.92 0.118
(0.303) (0.01)
$\log(n)\times$Size ($\beta_{1}$) $-0.178$
Table A2: Parameter estimates and standard errors ($K=5$).
Within Between
Estimate Estimate
(Std. Error) (Std. Error)
Edges ($\alpha_{1}$) $-4.893$ $-7.45$
(0.263) (0.006)
$\log(n)\times$ Edges $\left(\alpha_{1}^{\ell}\right)$ $-0.569$
Externality ($\gamma$) $-0.258$
Maturity ($\beta_{1}$) 0.711 $-0.267$
(0.302) (0.012)
$\log(n)\times$ Maturity $\left(\beta_{1}^{\ell}\right)$ $-0.012$
Popularity ($\beta_{1}$) 1.316 0.104
(0.31) (0.011)
$\log(n)\times$Popularity ($\beta_{1}$) $-0.111$
Size ($\beta_{1}$) 1.324 $-0.073$
(0.307) (0.011)
$\log(n)\times$Size $\left(\beta_{1}^{\ell}\right)$ $-0.142$
Table A3: Parameter estimates and standard errors ($K=15$).
Within Between
Estimate Estimate
(Std. Error) (Std. Error)
Edges ($\alpha_{1}$) $-4.181$ $-8.28$
(0.305) (0.006)
$\log(n)\times$ Edges $\left(\alpha_{1}^{\ell}\right)$ $-0.765$
Externality ($\gamma$) 0.135
Maturity ($\beta_{1}$) 3.201 $-0.198$
(0.321) (0.011)
$\log(n)\times$ Maturity $\left(\beta_{1}^{\ell}\right)$ $-0.307$
Popularity ($\beta_{1}$) 1.58 0.117
(0.287) (0.01)
$\log(n)\times$Popularity $\left(\beta_{1}^{\ell}\right)$ $-0.15$
Size ($\beta_{1}$) 0.369 $-0.074$
(0.29) (0.01)
$\log(n)\times$Size $\left(\beta_{1}^{\ell}\right)$ $-0.048$
ERGM Model
Estimate
(Std. Error))
Edges $-7.961$
$(0.007)$
Externality $-0.178$
$(0.002)$
Maturity $-0.131$
$(0.010)$
Size $0.127$
$(0.010)$
Popularity $-0.066$
$(0.010)$
Table A4: Parameter estimates and standard errors (K=1)

### v3

@@ Appendix A Exploiting sparse matrix operations for updating variational parameters
$\Lambda_{ik}$ $\left(\bm{g},\bm{x},\bm{\Xi}^{(t)}\right)=$ $\left(\sum_{\bm{x}\in\mathcal{X}}\sum_{d\in\{0,1\}}\bm{\Gamma}(d)\circ\bm{\Delta}(\bm{x})\,\bm{\Xi}^{(t)}\,\bm{\Pi}^{(t)}\left(d,\bm{x},\bm{\Xi}^{(t)}\right)+\left(\bm{\Gamma}(d)\circ\bm{\Delta}(\bm{x})\right)^{\top}\,\bm{\Xi}^{(t)}\,\bm{\Pi}^{(t)}\left(d,\bm{x},\bm{\Xi}^{(t)}\right)^{\top}\right)_{ik}.$
@@
$\bm{\pi}^{(t+1)}(1,\bm{x})=$ $\big{(}\left(\bm{\Xi}^{(t+1)}\right)^{\top}\bm{g}\circ\bm{X_{1}}^{x_{1}}\circ(\bm{J}-\bm{X_{1}})^{1-x_{1}}\circ\cdots\circ\bm{X_{p}}^{x_{p}}\circ(\bm{J}-\bm{X_{p}})^{1-x_{p}}\left(\bm{\Xi}^{(t+1)}\right)\big{)}\oslash$
$\left(\left(\bm{\Xi}^{(t+1)}\right)^{\top}\bm{X_{1}}^{x_{1}}\circ(\bm{J}-\bm{X_{1}})^{1-x_{1}}\circ\cdots\circ\bm{X_{p}}^{x_{p}}\circ(\bm{J}-\bm{X_{p}})^{1-x_{p}}\left(\bm{\Xi}^{(t+1)}\right)\right),$
@@
enabling the evaluation of $\bm{\Pi}\left(d,\bm{x},\bm{\Xi}^{(t)}\right)$ for $\bm{x}\in\mathcal{X}$. Following Dahbura et al. (2021), we can still evaluate this matrix by sparse matrix operations.
@@
In sum, we have shown how to compute the terms $\Omega_{ik}\left(\bm{0},\bm{0},\bm{\Xi}^{(t)}\right)$ and $\Lambda_{ik}\left(\bm{g},\bm{x},\bm{\Xi}^{(t)}\right)$ through matrix multiplications. Plugging in these terms into (A) enables a fast computation of $\Omega_{ik}(\bm{g},\bm{x},\bm{\Xi}^{(t)})$, which is the computational bottleneck for evaluating the quadratic term needed to update $\bm{\Xi}^{(t)}$ to $\bm{\Xi}^{(t+1)}$.
@@
Appendix B Additional figures and estimates: Comparison with other models
@@
In Figure A1 (A), we show the simulated statistics generated by a model with $K=1$, corresponding to an Exponential Random Graph Model (ERGM). In aggregate, this model performs very poorly, which suggests that the unobserved heterogeneity helps the model fit.
Figure A1: Comparison of model fits of network statistics of competing models.
@@
We compare our model with $K=10$ unobserved types to a stochastic blockmodel with $K=10$ unobserved types in Figure A1 (B). In essence, the comparison model is the same as our model without the externality. As expected, the stochastic blockmodel (SBM) performs worse in predicting the externality within blocks.
@@
We also performed robustness checks to choose the number of types $K$. In Figures A1 (C) and (D), we report the boxplot for the fit of models with $K=5$ and $K=15$, respectively.
@@
We conclude from the plots that increasing or decreasing the number of types does not seem to help in representing the externality. We report the estimates of all these competing models in Table A1.
@@
  ERGM (K = 1)   SBM (K = 10)
Within Between Within Between
Estimate Estimate Estimate Estimate
(Std. Error) (Std. Error) (Std. Error) (Std. Error)
Edges ($\alpha_{1}$) --7.961 --5.703 -7.979
(0.007) (0.274) (0.006)
@@
Externality ($\gamma$) --0.178
Maturity ($\beta_{1}$) --0.131 1.767 -0.222
(0.01) (0.303) (0.011)
$\log(n)\times$ Maturity ($\beta_{1}^{\ell}$) --0.12
Popularity ($\beta_{2}$) --0.066 1.071 -0.07
(0.01) (0.308) (0.011)
$\log(n)\times$Popularity ($\beta_{2}^{\ell}$) --0.117
Size ($\beta_{3}$) 0.127 1.92 0.118
(0.01) (0.303) (0.01)
$\log(n)\times$Size ($\beta_{3}^{\ell}$) --0.178
Table A1 (ctd.)   Local Dep. ERGM (K = 5)   Local Dep. ERGM (K = 15) Within Between Within Between Estimate Estimate Estimate Estimate (Std. Error) (Std. Error) (Std. Error) (Std. Error) Edges ($\alpha_{1}$) --4.893 -7.45 --4.181 -8.28 (0.263) (0.006) (0.305) (0.006) $\log(n)\times$ Edges ($\alpha_{2})$-0.569 --0.765 (0.028) (0.038) Externality ($\gamma$) --0.258 0.135 (0.013) (0.013) Maturity ($\beta_{1}$) 0.711 -0.267 3.201 -0.198 (0.302) (0.012) (0.321) (0.011) $\log(n)\times$ Maturity ($\beta_{1}^{\ell}$) --0.012 --0.307 (0.033) (0.041) Popularity ($\beta_{2}$) 1.324 -0.073 0.369 -0.074 (0.307) (0.011) (0.29) (0.01) $\log(n)\times$Popularity ($\beta_{2}^{\ell}$) --0.142 --0.048 (0.033) (0.037) Size ($\beta_{3}$) 1.316 0.104 1.58 0.117 (0.31) (0.011) (0.287) (0.01) $\log(n)\times$Size ($\beta_{3}^{\ell}$) --0.111 --0.15 (0.034) (0.037)